Solve
0.7(3x+4) greater than equal to 1.3-(x+2)

Answer by checkley77(12844)

0.7(3x+4)>=1.3-(x+2)
2.1x+2.8>=1.3-x-2
2.1x+x>=1.3-2-2.8
3.1x>=-3.5
x>=-3.5/3.1 answer.